Week of 160829

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Jesus (T0 Storage), Maarten (ALICE), Gavin (T0 Compute), Vincent (WLCG Security), Maria A. (WLCG Ops Coord and CMS), Julia (WLCG Ops Coord), Michal (ATLAS), Alberto (Monitoring), Daniele (CMS), Charles (T0 Databases), Luca (LHCb), Maria D. (WLCG Ops Coord and SCOD).
  • remote: Elena and Lucia (CNAF), Ulf (NDGF), David (FNAL), Rolf (IN2P3), Di (Triumf), Eric (BNL), Dennis (NL_T1), Kyle (OSG).

Experiments round table:

  • ATLAS reports (raw view) -
    • Activities:
      • normal running - grid is full
    • Problems:
      • RRC-KI-T1 finished the downtime - some files seemed to be unavailable (GGUS:123587) - files declared lost

Accesses and dumps.. then back since weekend to stable beams with 2220x2220 BCMS. Expected beam dump by LHC at 15:00 today.

  • SAM tests outage (Aug-25):
    • solved after a Condor restart, so the "real" SAM tests (=prod) are OK since Aug-25 at 5pm CERN time (not in preprod though)

  • CMS critical services, lxbatch (Aug-23): service unavailable for >1h
    • ELOG + service-now ticket submitted. OK after some hrs.

  • Queuing data traffic from T1_US_FNAL to PIC and JINR (Aug-23) -> GGUS:123559 (solved already)
    • persisted over several shifts. Debugged to be related to O(15) files on a problematic disk at the source site. Now OK.

  • Condor submit errors at T1_RU_JINR (Aug-27/28) -> GGUS:123608
    • seen as SAM3 CE errors, ELOG opened, then GGUS (solved with no explanation - to be checked)

  • !PhEDEx agents down at T1_RU_JINR (Aug-28) -> GGUS:123614
    • root cause: loss of network connection to Moscow.
    • not sure if we got the corresponding unscheduled downtime correctly, as it did not appear at the time of ticketing in either the CMS Site Downtime Calendar or in the SSB. To be checked.

Unfortunately Victor tried to connect from JINR but couldn't so there were no comments during the meeting as the problem was experienced already at 6am, i.e. before the dowtime recording in GOCDB

  • ALICE -
    • NTR

  • LHCb reports (raw view) -
    • Activity
      • Monte Carlo simulation, data reconstruction and user jobs on the Grid
    • Site Issues

Sites / Services round table:

  • ASGC: not connected
  • BNL: NTR
  • CNAF: A GGUS ticket about a data transfer issue is waiting for user.
  • EGI: NTR
  • FNAL: A scheduled downtime for Saturday 17/9 will start by shutting services down in the evening of 16/9. All expected to be back in the evening of 17/9.
  • GridPP: not connected
  • IN2P3: NTR
  • JINR: Batch stuck again at 27-08-2016, this case wasn't caught by watchdog. Lost both main and backup network links to Moscow at 28-08-2016, optics were cutoff, now back again.
  • KISTI: not connected
  • KIT: not connected
  • NDGF: Problem with ATLAS Storage SAM tests observed last Friday. Discovered wrong time (by 11'') on one GridFTP server.
  • NL-T1: NTR
  • NRC-KI: not connected
  • OSG: NTR
  • PIC: not connected
  • RAL: RAL Closed today and tomorrow (29/30 August)
  • TRIUMF: NTR

  • CERN computing services:
    • LCG router reboot on Thursday 1 September between 10.00 and 11.00 .. two ~7 minute outages. A GOCDB downtime (will be) declared. Will affect:
      • VOMS service
      • Lxbatch - some nodes will lose connectivity.
      • CEs ce403, ce404, ce406, ce407 and cesatltzero1.
      • Some VMs will also lose connectivity during this period (you will have been mailed details from network team).
Maybe the T0 should look into spreading some of these important services (in particular VOMS) over multiple routers.

  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring: NTR
  • MW Officer: not present
  • Networks: not present
  • Security: NTR

AOB:

  • Ops Coord meeting this Thu 1/9 at 15h30 CEST with theme CentOS7. Agenda http://indico.cern.ch/event/540420/
  • Dennis reported that vidyo crashed a few secs after upgrade. Maria D. as SCOD will open a SNOW ticket.
Edit | Attach | Watch | Print version | History: r10 < r9 < r8 < r7 < r6 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r10 - 2016-08-29 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback